### v1

# iPanda: An Intelligent Protocol Testing and Debugging Agent for Conformance Testing

Abstract.
Conformance testing is essential for ensuring that protocol implementations comply with their specifications. However, traditional testing approaches involve manually creating numerous test cases and scripts, making the process labor-intensive and inefficient. Recently, Large Language Models (LLMs) have demonstrated impressive text comprehension and code generation abilities, providing promising opportunities for automation. In this paper, we propose iPanda, the first end-to-end framework that leverages LLMs to automate protocol conformance testing. Given a protocol specification document and its implementation, iPanda first employs a keyword-based method to automatically generate comprehensive test cases. Then, it utilizes a code-based retrieval-augmented generation approach to effectively interpret the implementation and produce executable test code. To further enhance code quality, iPanda incorporates an iterative self-correction mechanism to refine generated test scripts interactively. Finally, by executing and analyzing the generated tests, iPanda systematically verifies compliance between implementations and protocol specifications. Comprehensive experiments on various protocols show that iPanda significantly outperforms pure LLM-based approaches, improving the success rate () of test-code generation by factors ranging from to .
1. Introduction
Communication protocols such as HTTP, MQTT, and CoAP form the backbone of today’s information-driven society, playing critical roles in internet data transmission, IoT connectivity, and cloud computing services. Ensuring the stability and efficiency of protocol implementations is therefore paramount and necessitates rigorous testing procedures. Conformance testing, specifically designed to verify adherence of protocol implementations to their official standards and specifications, is one of the essential methods employed in validating communication protocols. In typical scenarios, testers might spend weeks manually writing extensive test scripts utilizing protocol implementation libraries, and individually verifying compliance with protocol requirements. For instance, a typical compliance check of MQTT protocol implementations might require engineers to manually develop hundreds of script-based tests, each verifying different message formats and states. With the increasing complexity of protocols, traditional manual testing methods have become inefficient, cumbersome, and difficult to generalize. This situation underscores an urgent need for automated testing solutions that minimize human intervention while enhancing efficiency and coverage.
Concurrently, large language models (LLMs) have shown impressive capabilities in language comprehension, reasoning, and performing sophisticated tasks. Trained on extensive datasets, LLMs effectively generalize to novel tasks, parse complex inputs, and maintain context, making them ideal for automating intricate operations. For instance, in IoT fuzz testing, LLM-driven automation significantly enhances protocol message generation, increasing vulnerability detection effectiveness and uncovering previously undetected issues (Wang et al., 2024). Similarly, in mobile device automation, LLMs combine general reasoning with domain-specific expertise, facilitating complex tasks without extensive manual scripting (Wen et al., 2024a, b). These examples demonstrate the transformative potential of LLMs in turning labor-intensive procedures into efficient automated solutions.
Given this context, a compelling question arises: Can we leverage the capabilities of LLMs to address the challenges inherent in conformance testing of communication protocols? However, the practical realization of this vision faces several significant hurdles:
Efficiently generating comprehensive test cases that rigorously satisfy protocol testing requirements remains a significant and challenging task.
Equipping LLMs with the capability to accurately interpret, adapt, and interact with existing protocol implementation libraries poses considerable difficulty, due to varying specifications and complexity.
Improving the overall performance of LLM-driven testing processes is vital, especially regarding the accuracy, effectiveness, and reliability of the generated test code, ensuring conformance and dependability in real-world testing.
Addressing these challenges is essential not only to resolve the immediate practical issues but also to establish foundational methodologies as LLM technologies continue to advance in network protocol testing.
Motivated by trust in the adaptive reasoning capabilities of LLMs and inspired by successful applications in related domains, we tackle the aforementioned challenges by introducing iPanda, an Intelligent Protocol Testing and Debugging Agent. To the best of our knowledge, iPanda represents the first end-to-end, LLM-based solution specifically designed for automating communication protocol conformance testing. It streamlines the entire testing process—from protocol specification to result analysis—significantly reducing manual effort and improving overall efficiency.
Specifically, iPanda leverages large language models to autonomously generate comprehensive test cases directly from protocol specification documents. Given a specific protocol implementation library, iPanda dynamically generates executable test-case code, runs these tests within a simulated environment, and systematically identifies potential compliance issues. To efficiently generate test cases aligned with precise testing requirements, we propose a novel keyword-based test-case generation approach grounded in the inherent characteristics of protocol specifications.
Moreover, to ensure the rapid adaptation of iPanda’s core (i.e., the underlying LLMs) to varying protocol libraries, we developed a code-based retrieval-augmented generation (RAG) mechanism. This mechanism guides the agent dynamically and accurately in employing diverse protocol implementation libraries. Additionally, to further enhance the accuracy and reliability of the generated test code, we incorporate a code self-correction mechanism. This allows the LLM to iteratively validate and refine its outputs interactively, closely mimicking human debugging procedures.
iPanda also supports natural-language commands and demonstrates reasoning capabilities for complex state transitions, making it particularly effective for dynamic, heterogeneous network environments where traditional testing methods frequently fall short. By significantly reducing the dependency on human expertise and fully automating conformance test execution, iPanda presents a transformative advancement in communication protocol conformance testing.
Our contributions are as follows:
To the best of our knowledge, we present the first automated protocol conformance testing framework integrating LLMs with domain-specific expertise, including keyword-based test case generation, code-based RAG, and a self-correction mechanism for test-code refinement.
We design and implement iPanda, an intelligent, LLM-powered agent capable of autonomously extracting test cases from protocol specifications, invoking protocol implementation libraries, executing and debugging tests, and identifying conformance issues.
Comprehensive experiments demonstrate that, compared to the pure LLM-based method, iPanda improves the of test code generation by to .
Upon acceptance of this paper, we will publicly release the associated source code and datasets to facilitate future research and practical applications.
2. Background and motivation
2.1. Conformance testing of protocols
Conformance testing for communication protocols is a method used to verify whether a protocol implementation complies with the requirements defined by the protocol specification. Its primary goal is to ensure that the protocol implementation behaves in a conformant manner across various scenarios and input conditions, thereby guaranteeing compatibility and interoperability between different protocol implementations. A complete protocol conformance testing process typically includes the following key steps:
Specification analysis. Thoroughly understand and analyze the protocol’s standard documents to extract explicitly defined behaviors and requirements.
Test case design. Develop comprehensive test cases based on the protocol specifications.
Test case implementation. Write test scripts or code using the protocol implementation under test to concretely realize the test cases.
Test execution. Run the test case code in a properly configured testing environment and collect results.
Result analysis and evaluation. Analyze the test results to determine whether the system under test complies with the specified requirements.
2.2. Large language models
Large Language Models (LLMs) are artificial intelligence models built on deep learning techniques, pre-trained on vast amounts of textual data to capture rich semantic and contextual relationships in language. Prominent LLMs, such as the GPT series (Jaech et al., 2024; Achiam et al., 2023; Hurst et al., 2024), LLaMA series (Touvron et al., 2023; Grattafiori et al., 2024), and DeepSeek series (Guo et al., 2025; Liu et al., 2024), possess the capability to generate detailed content, understand complex semantics, and perform a wide range of tasks, including text generation, question answering, and code writing.
To better leverage LLM capabilities, researchers introduced prompt engineering, carefully designing prompts or instructions to guide models toward generating more accurate and effective outputs. Building upon this, a series of prompt augmentation techniques have emerged, further improving model adaptability and generalization. This work specifically adopts several such methods: Chain of Thought (CoT) (Wei et al., 2022), instructing models to explicitly provide intermediate reasoning steps to better handle complex reasoning tasks; In-Context Learning (ICL) (Dong et al., 2022), guiding models with example-based formats to rapidly adapt without additional training; Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), integrating external knowledge or retrieval mechanisms to enhance performance on knowledge-intensive tasks; and Self-Correcting with Tool-Interactive Critiquing (CRITIC) (Alinezhad et al., 2019), iteratively refining outputs through external tool-driven feedback. Together, these approaches significantly boost LLM effectiveness and adaptability in real-world applications.
2.3. Integrating LLM and conformance testing
The current conformance testing process has significant limitations, primarily due to its heavy reliance on manual effort. Conformance testing involves a vast number of detailed test cases that must be manually written, making the process labor-intensive and time-consuming. Additionally, implementing these test cases requires extensive test code development, further increasing the cost of development and maintenance. These limitations significantly reduce the efficiency and flexibility of conformance testing.
At present, no end-to-end conformance testing tool exists that seamlessly integrates protocol documents, implementations, and test results. The exceptional text comprehension and code generation capabilities of LLMs offer new possibilities for addressing these shortcomings. By leveraging the powerful language understanding and generation abilities of LLMs, test case design and test code development can be automated or semi-automated, substantially reducing the manual workload. Therefore, this work aims to integrate LLMs into the conformance testing process for communication protocols. We propose an end-to-end conformance testing agent framework to enhance the efficiency and effectiveness of conformance testing.
3. Design of iPanda
In this section, we provide a detailed introduction to the design of iPanda. This system is specifically designed for communication protocol conformance testing, particularly in network environments such as the Internet of Things. iPanda can dynamically generate test case sets for conformance testing based on protocol specification documents. For a given protocol implementation library, it can generate test case code, interact with the implementation in a simulated communication environment, perform testing and debugging, and analyze execution results to identify potential deficiencies in the implementation.
3.1. Overview
The overview of iPanda is shown in Fig. 1. Taking the RSocket protocol (rsocket, 2024) as an example, suppose a user wants to verify whether its Python implementation rsocket-py (rsocket py, 2025) conforms to its specification. iPanda first extracts key functional points from the protocol document and automatically generates standardized test cases using the LLM generator, optionally applying a filter to remove anomalous cases. For each test case, iPanda generates executable test code using the target implementation library. It retrieves relevant context from the implementation library, integrating this context into detailed prompts that clearly define the LLM’s role and task objectives. To ensure executability, the generated code undergoes validation; if issues are detected, iPanda initiates iterative refinement using historical context and error information, dynamically adjusting prompts to debug the code. Once the generation yields executable code or reaches a retry limit, iPanda compiles a final debugging report and evaluates test outcomes, determining compliance with the protocol’s conformance requirements.
To address the challenges mentioned in Sec. 1 and enhance iPanda’s effectiveness in test cases and code generation, we have designed several optimization methods within iPanda. These techniques are embedded within individual modules or integrated across multiple modules. The following sub-sections will provide a detailed explanation of these optimizations.
3.2. Test case generation
3.2.1. Test case generation methods
Traditional test case generation methods include:
Specification-driven methods. Based on the protocol’s standard specification, these methods analyze state machines and message interaction flows to design test cases that cover both valid and invalid scenarios.
Model-driven methods. By constructing formal models such as finite state machines and extended finite state machines, these methods automatically generate test sequences to ensure comprehensive coverage of states and message transitions.
Implementation-driven methods. These methods test the completeness and robustness of a protocol implementation by analyzing code coverage and injecting faults based on the specific implementation.
Interoperability testing methods. These methods aim at verifying compatibility between different vendors and versions of a protocol, ensuring correct communication between systems.
Random and fuzz testing methods. These methods use random messages and anomalous data inputs to uncover hidden defects in the protocol implementation, improving security and fault tolerance.
The above methods often require significant human and material resources from professionals, lacking sufficient flexibility and scalability. For example, in the model-driven methods, constructing a formal model specific to a communication protocol is necessary, but this model cannot be easily extended to other protocol conformance testing tasks.
Given LLMs’ great text comprehension and generation capabilities, they are naturally suited for analyzing and extracting information from specification documents. For instance, even without fine-tuning, an LLM can identify key statements from a given RFC document simply by being prompted to extract relevant textual information. However, due to the inherent randomness of LLMs, they do not strictly adhere to instructions, making extracted information potentially nonconformant or uncontrolled. Therefore, solely relying on LLMs for the direct extraction of key information is impractical and unreliable.
3.2.2. Keyword-based test case generation method
Insight: The majority of protocol documents follow strict format requirements and are structured using specific terminology to define and organize key concepts.
Taking CoAP as an example, we reviewed a total of RFC documents related to CoAP and found that of them comply with RFC 2119 (Shelby et al., 2014; Editor, 2025). These documents use uppercase keywords such as MUST, MUST NOT, REQUIRED, SHALL, and SHALL NOT, to define and emphasize critical protocol specifications and requirements. Even those documents that do not strictly follow RFC 2119 still use similar auxiliary verbs to highlight protocol constraints.
Inspired by this insight, we design a novel specification-driven test case generation method named keyword-based test case generation (keyword-based TCG). This method is integrated into the test case generation module, as shown in Fig. 1. Keyword-based TCG combines heuristic rules with the idea of generating datasets using LLMs. Specifically, this method first detects whether the document follows RFC 2119. If it does, keyword-based TCG uses regular expressions to extract paragraphs containing uppercase keywords. Each extracted paragraph is a complete natural paragraph to preserve as much contextual semantic information as possible. We define these paragraphs as functional points. If the document does not conform to RFC 2119, the method defaults to case-insensitive keyword extraction and repeats the above process.
After extracting functional points, keyword-based TCG calls the LLM to generate test cases based on these points. To ensure a standardized format for the generated test cases, we introduce few-shot in-context learning (few-shot ICL) into the LLM. Specifically, we construct a prompt with input-output example , where represents a functional point, and is a test case in the desired format. These input-output pairs are concatenated in the format:. During reasoning, the test is appended to the prompt, and the LLM learns the structure from the provided examples, generating an output in the same format, as illustrated in Fig. 2. To ensure the generated test cases are both domain-relevant and effective, we introduce a filter, allowing users to review the functional points and their corresponding generated test cases, removing any anomalous test cases.
Remark: It is important to note that the test case generation module is optional, meaning that test cases do not necessarily have to be generated by this module. iPanda allows users to import their own test cases, meanwhile still providing compatibility with subsequent code generation and conformance verification functionalities. The main purposes of the test case generation module are twofold: to facilitate the rapid generation of effective test cases directly from protocol documents, reducing the workload for users; and to provide standardized experimental datasets for this work, enabling a quantitative evaluation of iPanda’s performance.
3.3. Code-oriented prompting
3.3.1. Potential issues in LLMs
Although LLMs have demonstrated exceptional performance, their outputs are not entirely reliable. Taking code generation as an example, even when restricting the scope to a certain protocol implementation library, LLMs still exhibit the following issues during the code generation process:
LLM’s limited understanding of code libraries. We analyzed the cutoff dates of training data for several mainstream base LLMs, as shown in Tab. 1. Due to the delay in data acquisition, even the most recent base LLMs are trained on data that is only up to date as of a few months prior. For new implementation libraries, if they are not promptly included in the training dataset, these implementations remain unknown to the LLM. Furthermore, a scarcity of open-source examples for implementation libraries also results in the lack of training data, making it difficult for LLMs to efficiently generate code based on these libraries. This phenomenon will be further demonstrated in subsequent experiments.
Hallucination issues in LLM-generated code. The hallucination problem refers to instances where the LLM generates outputs that appear realistic and credible but are actually incorrect, fabricated, or baseless. This issue stems from the probabilistic prediction mechanisms inherent in statistical learning and is difficult to eliminate completely. In this work, hallucinations in LLM-generated code include calling non-existent classes or attributes from the implementation library, incorrectly using classes or methods in the implementation library, and incorrectly configuring the required parameters.
These issues lead to inefficiencies when blindly relying on LLMs to generate task-specific code based on a given implementation library.
| LLM | Release date | cutoff |
| GPT-4.5-preview (OpenAI, 2025) | 2025/2/27 | 2023/10 |
| GPT-4o(2024-11-20) (Hurst et al., 2024) | 2024/11/20 | 2023/10 |
| Deepseek-R1 (Guo et al., 2025) | 2025/1/20 | unknown |
| Deepseek-V3 (Liu et al., 2024) | 2024/12/26 | unknown |
| Claude-3-7-sonnet (Anthropic, 2025) | 2025/2/19 | 2024/11 |
| Gemini-2.0-flash (Google, 2025) | 2025/2/5 | 2024/1 |
| Qwen2.5-Coder-32B (Hui et al., 2024) | 2024/11/6 | unknown |
3.3.2. codeRAG
Insight: For newly developed protocol implementation libraries, the code details and examples in these libraries naturally contain high-quality, protocol-specific knowledge that LLMs may lack.
LLMs can efficiently learn and utilize new knowledge through in-context learning. This forms the foundation for the effectiveness of Retrieval-Augmented Generation (RAG). To address the issues mentioned above, inspired by the above insight, we design a code-based RAG mechanism named codeRAG, which dynamically guides iPanda in utilizing protocol implementation libraries. As shown in Fig. 3, given a protocol implementation library consisting of source code documents , we first use an embedding model to map these documents into their corresponding embedding representations . To enhance the prompt at the current step , we apply the same embedding model to map it into the same feature space, obtaining its corresponding embedding . The retriever then evaluates the similarity between and other embeddings using the cosine similarity formula:
where the larger the similarity score , the more relevant the document is to the current prompt . By default, we select the four source code documents with the highest similarity scores as reference examples and append them to the prompt for augment. This allows the LLM to leverage few-shot in-context learning to learn how to use the protocol implementation library, leading to more precise code generation.
It is important to note that codeRAG plays a role in both the startup stage (generating code for the first time) and iteration stage (iteratively correcting the code) of iPanda’s code generation process. In the startup stage, the retrieved source code documents help enhance the LLM’s understanding of the protocol implementation. In the subsequent iteration stage, these documents assist in mitigating the LLM’s hallucination issues.
We illustrate the effectiveness of codeRAG with an example. Taking rsocket-py as a case study, when generating code directly from test cases, the LLM may call non-existent methods in the classes due to its limited understanding of rsocket-py. By incorporating the retrieved source code of the relevant class, the LLM can self-correct these bugs. Furthermore, subsequent experiments in Sec. 4.4 also confirm the effectiveness of introducing codeRAG.
As part of codeRAG, the source code of protocol implementation, embedding model, and retriever are integrated into the long-term memory repository of the memory module. At the same time, the long-term memory repository also stores executable code documents generated from previous tests and historical debugging experience summarized by the summarizing module. In addition to retrieving the source code, codeRAG also retrieves these stored knowledge assets, further improving iPanda’s capability in automatic code synthesis and debugging.
3.3.3. Augmented Role Prompting
Prompt engineering is one of the key techniques for improving the quality of LLM’s outputs (White et al., 2023; Giray, 2023; Ekin, 2023). By carefully designing prompts, we can guide the LLM to generate results that better align with expectations. In Sec. 3.3.2, we introduced the optimization technique of using RAG. However, the complete structure of the final prompt used for LLM inference remains unclear. In this sub-section, we present the Augmented Role Prompting method used in iPanda.
In iPanda, we adopt and augment the Role Prompting method to generate prompt templates. Role Prompting explicitly assigns a specific role to the LLM within the prompt, guiding it to perform tasks more effectively, and improving response conformance, accuracy, and professionalism. Generally, a basic Role Prompting structure includes the following components:
Role. It specifies the particular role that the LLM should assume.
Task. It clearly defines the task that the LLM needs to perform. In the startup stage, the task is to generate code, while in the iteration stage, the task is debugging.
Context. It represents the contextual information that the LLM may utilize, i.e., the knowledge retrieved using the codeRAG method from the long-term memory repository, as discussed in Sec. 3.3.2.
In addition to the components mentioned above, the Augmented Role Prompting also introduces instruction to control the LLM’s output. The instruction component primarily includes the following elements:
Zero-shot CoT Prompting. It activates the LLM’s reasoning capabilities to improve the quality of code generation.
Additional guiding statements. They encourage the LLM to actively utilize context knowledge, automate decision-making in code generation, and facilitate debugging.
LLM output requirements. Specifically, the LLM generates outputs in JSON format based on a predefined output template. The JSON output template is structured as follows:
"script1": "print(’Hello, script1’)",
"script2": "print(’Hello, script2’)",
"script3": "print(’Hello, script3’)",
"execution_order": [
["script2", "script3"],
["script1"] ]
, where execution_order controls the execution order of the generated code. The scripts belonging to the same list are executed in parallel.
The Augmented Role Prompting is integrated into the prompt generator of the reasoning module. The complete generated prompt is shown in Fig. 4, where represents the prompt used in the startup stage, while denotes the prompt used in the iteration stage.
3.4. Optimizing code generation
3.4.1. Shortcomings in the startup stage
During the startup stage of iPanda, although the LLM relies on codeRAG and Augmented Role Prompting, effectively reducing the likelihood of hallucinations, we still cannot guarantee that the test code generated by the LLM is both accurate and executable. This issue arises from the inherent limitations of LLMs, which have been thoroughly discussed in Sec. 3.3.1. Therefore, it is necessary to introduce more sophisticated mechanisms to improve the quality of code generation.
3.4.2. Iteration stage: augmented CRITIC
Insight: For code generation, error messages encountered during code execution serve as a naturally accessible and high-quality source of knowledge.
Naturally, the most straightforward way to verify the quality of generated code is to execute it and observe the results. Whether it is the output from a successful execution or a bug report from a failed execution, this information provides additional insights for the LLM. Such feedback can effectively guide the LLM in refining and debugging.
Inspired by this insight, we introduce the iteration stage. The reasoning module’s LLM utilizes an external code execution tool from the execution module to iteratively validate and refine its output through human-like interaction. This process continues until either executable code is generated or the maximum number of retries is reached. This method, known as Self-Correcting with Tool-Interactive Critiquing (CRITIC), was initially proposed in (Alinezhad et al., 2019). Specifically, given the reasoning module’s LLM model and the input test case , the initial code solution is generated by the prompt as . The execution module’s tool then tests this initial solution, producing critiques . For the -th iteration (), the LLM’s output iterates as , where . In the context of code generation, at -th step consists only of the code generated at the current step. Thus, the iterative process does not retain a debugging trajectory of previous code versions. To address this, we augment CRITIC by modifying the LLM’s output at step as:
where is a prompt template designed to guide the LLM in debugging. The parameter represents the window size of the short-term memory cache, preventing the input length from exceeding the maximum text length that the LLM can handle (e.g., GPT-4o supports up to 128K tokens (Hurst et al., 2024)). When , the augmented CRITIC degrades to a codeRAG-based CRITIC; when , it considers all historical interaction information. The is the context retrieved using the codeRAG retriever , specifically as:
For code generation, the critiques returned by the execution module play a crucial role in the debugging process, as they identify code errors and provide actionable modification suggestions. These critiques guide the next iteration of code, helping to avoid similar mistakes. Inspired by the human iterative drafting and refinement process, we adopt an iterative validation-correction method. This process continues until the generated code meets specific execution success criteria or the maximum step of iterations is reached. The pseudocode for augmented CRITIC is shown in Alg. 1.
The augmented CRITIC is integrated into the reasoning module, execution module, and memory module. It operates through textual interactions among these modules to achieve iterative refinement. Specifically, the interaction between the reasoning module and the execution module is structured as a JSON object containing the test code and execution sequence, following the JSON template mentioned before. The interaction between the execution module and the memory module also involves JSON objects, which record the execution results or error reports for each generated code. Meanwhile, the interaction between the reasoning module and the memory module is text-based, embedding both the historical JSON strings cached in the short-term memory cache and the code retrieved from the long-term memory repository using the codeRAG.
Remark: During the iteration stage of iPanda, the test code is continuously refined. As the number of iterations increases, the length of the input context grows linearly, which in turn increases the reasoning load on the LLM. Therefore, it is crucial to set an appropriate cache window size to limit the context length. Additionally, as the number of iterations increases, the marginal benefit of code correction diminishes. Hence, selecting an optimal maximum step of iterations is essential. Based on experimental observations, we set the default maximum step to .
3.5. Summarizing bugs
The summarization module performs two key functions:
Summarizing code generation experience. Under appropriate prompt guidance, the summarization process generates experience by analyzing the code testing process , where does not exceed the predefined maximum step of LLM iterative reasoning. The experience is stored in the long-term memory repository as valuable knowledge to continuously enhance iPanda’s overall performance. By leveraging the accumulated experience, the augmented CRITIC can refine code corrections more efficiently, ensuring that iPanda continuously learns and improves its ability to handle similar tasks in the future with greater efficiency and accuracy.
Ensuring conformance between protocol and its implementation. With appropriate prompt guidance, the LLM also reviews the test cases and the usage of the protocol implementation library. It evaluates whether the tested protocol implementation adheres to the protocol specifications embedded in the test cases, and generates a corresponding conformance testing report.
Due to the inherent randomness in response generation and the LLM’s imperfect adherence to instructions, the LLM may mistakenly interpret the task of conformance tasting as summarizing bugs encountered during the code generation process. Based on experimental observations, the common types of bugs identified by the LLM in code generation primarily include incorrect parameter passing (e.g., incorrectly binding to any-address) and incorrect method or attribute calls (e.g., some attribute does not exist). To address this issue, a keyword-based filtering method can be applied to the conformance testing reports. This filtering process removes portions of the report that contain keywords related to code generation bugs. The remaining reports are then subjected to manual review and verification to ensure the accuracy of the conformance testing results.
4. Evaluation
In this section, we present our experimental setup, all experimental results and analyses, to demonstrate the effectiveness of iPanda.
4.1. Experimental setup
4.1.1. Experimental platform
Hardware. We deployed the iPanda prototype system on a local server, equipped with a 32-core 13th Gen Intel(R) Core(TM) i9-13900HX CPU and an NVIDIA GeForce RTX 4060 Laptop GPU. To maximize the potential of iPanda, we utilized a cloud-based LLM API for reasoning, meaning that the local device is primarily responsible for managing iPanda’s process control and long-term memory repository. As a result, the demand for CPU and GPU resources is relatively low. The GPU is only required to support the normal operation of the embedding model.
Software. We implemented iPanda using Python. All target protocol implementation libraries, along with their required libraries, were pre-installed in a local virtual environment. For LLM, we used GPT-4o (Hurst et al., 2024), DeepSeek-V3 (Liu et al., 2024), and Qwen2.5-Coder-32B (Hui et al., 2024). These models represent the most advanced base models, differing in architecture, size, and pretraining focus. Notably, we deliberately included a smaller-scale Coder model to examine how a code-specialized model, fine-tuned for coding tasks, adapts to the conformance testing task (Suzgun et al., 2022; Shao et al., 2024). All LLMs were configured with short-term memory window size , , and . However, since the LLM services were accessed via APIs, slight randomness remained in the generated responses. For codeRAG, we utilized OpenAI’s text-embedding-3-large as embedding model (OpenAI, 2024). For simplicity and generality, the testing environment was configured locally, using the local IP address and different ports to simulate network conditions.
4.1.2. Tested protocols and Implementations
We selected the following protocols and their corresponding Python implementation libraries as test subjects:
CoAP & aiocoap (aiocoap, 2025). The Constrained Application Protocol (CoAP) is a lightweight network communication protocol designed for resource-constrained devices, primarily used for IoT devices. It was standardized in 2014 as RFC 7252 (Shelby et al., 2014). As of March 2025, there are RFC documents related to CoAP (Editor, 2025). We selected aiocoap, the Python implementation of CoAP, as the target for conformance testing. As of March 2025, aiocoap fully or partially supports CoAP-related RFC documents.
RSocket & rsocket-py. RSocket is an application protocol that provides reactive stream semantics over an asynchronous, binary boundary. As of March 2025, RSocket has not yet been formalized as an RFC standard. However, it has a well-established protocol specification (rsocket, 2024) (adhering to RFC 2119) and multiple implementations across various programming languages. We selected rsocket-py, the Python implementation of RSocket, as the target for conformance testing.
The key reasons for selecting these two protocols and their implementation libraries for testing are that they meet important criteria for conformance testing:
The protocols have established mature standards and specifications, but their implementations have only recently developed.
Only partial specifications have been implemented in these protocol libraries.
There is a lack of mature conformance testing tools for these implementations.
It is important to note that LLMs may already contain embedded knowledge about the selected protocols and their implementation libraries. While the RAG is to supplement the LLM with missing knowledge. Therefore, through explicit interaction with the LLM, we inferred the following conclusions. For aiocoap, the LLM possesses extensive built-in knowledge about aiocoap. Consequently, when testing it, we froze the long-term memory repository to prevent unnecessary external influence. For rsocket-py, due to its relatively recent development, and limited open-source examples, the LLM lacks accurate knowledge of it. Therefore, when testing it, we retained the long-term memory repository and incorporated its open-source code. Subsequent experiments confirmed the validity of these operations.
4.1.3. Dataset
Using the keyword-based TCG in Sec. 3.2.2, we introduced two test case sets for CoAP and RSocket:
CoAP-set. We selected RFC documents directly related to CoAP to generate the CoAP-set, which consists of uniformly formatted test cases.
RSocket-set. We generated the RSocket-set using RSocket document, having test cases.
All test cases underwent professional manual review and selection to ensure their authenticity and validity. Each test case follows a standardized format comprising five components: test case name, test preconditions, test steps, test assertion, and precautions. It is important to note that the core focus of our work is on the accuracy of iPanda’s code generation and the effectiveness of its conformance testing. Therefore, we do not aim for the test case sets to comprehensively cover all protocol specifications. Instead, we ensure the rationality of the test cases and maintain conformance across all experiments by using the same set of test cases.
4.1.4. Baseline
Specification-driven conformance testing methods typically require the development of protocol-specific conformance testing tools. Unfortunately, no open-source conformance testing tools currently exist for CoAP and RSocket. Moreover, our work is the first to employ LLMs for conformance testing in a protocol-agnostic manner. To reasonably evaluate iPanda’s performance, we selected a pure-LLM baseline approach in which the LLM generates test code in a single step (using the same prompt and randomness parameter settings). This corresponds to the startup stage of iPanda with the memory module disabled, serving as our experimental baseline.
4.1.5. Metrics
. This metric is commonly used to evaluate the correctness and reliability in generating code for a given programming task, which is a core sub-task of iPanda. measures the probability that at least one of the generated code solutions is correct. Given the inherent randomness in LLM-generated outputs, we can simplify this metric. Specifically, for a test set containing cases , the metric for a LLM is defined as:
where denotes the XOR operator, and represents the correctness evaluation of the -th generated solution. It takes a value of if the solution is correct and otherwise.
Conformance testing results. This is a qualitative metric that includes both positive sample (generating executable code) and negative sample (generating unexecutable code due to library bugs possibly) analyses based on whether the protocol implementation complies with the specification requirements. It is used to demonstrate the effectiveness of iPanda in conformance testing.
4.2. Marginal benefit analysis in the iteration stage
As the number of iterations increases, the marginal benefit of code correction in the iteration stage gradually decreases. To achieve the optimal marginal benefit, we must first determine an appropriate maximum step of LLM iterative reasoning. In our experiments, we set the upper limit for reasoning iterations to . iPanda was tested on the CoAP-set, which contains a larger number of test samples. To analyze the required number of reasoning iterations for successfully generating executable code, we constructed a histogram depicting the cost of successful generations. The experimental results are shown in Fig. 5. It can be observed that in most cases, no more than steps are required to generate executable code. Specifically, when setting the maximum inference iterations to , using GPT-4o, iPanda successfully tested cases in the CoAP-set (classified as positive samples), among which cases required no more than iterations, accounting for . Using DeepSeek-V3, the corresponding proportion was . Using Qwen2.5-Coder-32B, the proportion was . Under the influence of the scaling law, large-scale foundation models such as GPT-4o and DeepSeek-V3 (both exceeding parameters) exhibit stable performance in code generation and correction. As the number of iterations increases, the marginal benefit steadily declines. However, for Qwen2.5-Coder-32B, which has only parameters, the marginal benefit is relatively unstable. This may be attributed to the model’s smaller scale, leading to weaker code generation capabilities. By default, we set the maximum step of LLM reasoning iterations to in subsequent experiments to achieve the optimal marginal benefit.
4.3. Performance analysis of code generation
We conducted experiments using the baseline and iPanda supported by the three LLMs to evaluate their code generation performance on the CoAP-set. The experimental results for all methods are shown in Tab. 2. Experimental results show that under the same LLM conditions, iPanda significantly enhances code generation capabilities compared to the baseline. is improved by (GPT-4o), (Deepseek-V3), and (Qwen2.5-Coder-32B) respectively. For example, when using the GPT-4o model, the baseline achieves only a of . In contrast, iPanda improves this by , reaching . This strongly validates that aiocoap aligns with the specifications underlying these test cases, reducing the likelihood of nonconformances and narrowing the focus for critical testing. Even with the smallest model Qwen2.5-Coder-32B, iPanda’s is nearly twice that of the baseline using the most advanced model GPT-4o. These results confirm the success of the augmented CRITIC method. With its support, even smaller-scale LLMs can be stimulated to exhibit stronger code generation capabilities, achieving performance comparable to or even surpassing larger, more advanced LLMs.
| Used-LLM | Baseline | iPanda |
| GPT-4o | 17.32% | 80.95% |
| Deepseek-V3 | 9.52% | 57.14% |
| Qwen2.5-Coder-32B | 3.03% | 32.03% |
Due to the inherent randomness in LLM-generated content, multiple repeated tests are typically required for code generation, and performance is often evaluated using . Therefore, we conducted a joint experiment to examine the impact of the maximum reasoning iterations and the number of repeated tests . Given the strong performance of GPT-4o, we selected it as the default LLM for iPanda. The experiment was configured with and . The experimental results on the CoAP-set are shown in Fig. 6-(a), where the numbers in squares represent the number of executable code generated by iPanda under the current and . When setting and , the experiment degenerates into the setup of Fig. 5-(a). When setting and , the experiment degenerates into the baseline evaluation. From the results, we observe that increasing the reasoning iterations significantly improves iPanda’s performance. Additionally, repeated testing can also enhance the success rate of code generation, though its impact is considerably smaller than increasing reasoning iterations. For instance, increasing the number of repeated tests by five (i.e., instead of ) raises the number of positive samples from to . Increasing the number of inference iterations by five (i.e., instead of ) raises the number of positive samples from to . These results validate the rationality of our decision to adopt a sequential strategy (i.e., iterative code generation using a single LLM), rather than a parallel strategy (i.e., involving multiple LLMs generating code simultaneously). While the former has a higher interaction cost per step due to increased context length, this additional context contains valuable execution feedback, which provides richer information to guide the LLM in refining the generated code.
4.4. Ablation study
To validate the effectiveness of the methods used in iPanda, we conducted an ablation study to analyze the contributions of codeRAG and augmented CRITIC. The experimental results are presented in Tab. 3.
| Approach | CoAP-set | RSocket-set |
| iPanda | 80.95% | 38.71% |
| iPanda, no codeRAG | 80.95% | 14.51% |
| iPanda, no augmented CRITIC | 17.32% | 3.23% |
| Baseline | 17.32% | 11.29% |
Since LLMs have a deep understanding of CoAP and aiocoap, the long-term memory repository was frozen, effectively disabling the codeRAG mechanism. As a result, codeRAG does not impact the testing outcomes for this protocol, leading to only two experimental configurations in the ablation study. To evaluate the effectiveness of codeRAG, we introduced additional tests for the RSocket protocol and its Python implementation, rsocket-py. In this setup, iPanda was preloaded with the open-source code of rsocket-py for use in codeRAG. Experimental results on the RSocket-set revealed that the inclusion of codeRAG significantly improved LLM’s ability to understand and correctly utilize rsocket-py. This resulted in an increase in from to .
When the augmented CRITIC is removed, iPanda degrades into a baseline method that only incorporates codeRAG. Experimental results show that, on both datasets, the performance of iPanda drops significantly without the augmented CRITIC. This phenomenon has been extensively studied in previous experiments. Nevertheless, on the RSocket-set, even without the augmented CRITIC, iPanda still outperforms the baseline method. This further demonstrates the effectiveness of codeRAG.
Remark: The primary design goal of iPanda is to support conformance testing for multiple communication protocols. Therefore, it must exhibit protocol compatibility. At the same time, we also expect iPanda’s performance to improve as the underlying LLMs advance, requiring it to maintain model compatibility. The extensive performance and ablation experiments conducted above confirm that iPanda demonstrates strong compatibility with both different protocols and different LLMs.
4.5. Results analysis of conformance testing
According to the joint experiment results in Sec. 4.3, maximizing the number of iterations and the number of repetitions yields the highest number of positive samples, indicating that aiocoap conforms to the majority of protocol specifications. For negative samples, we analyzed the documents they belong to, as shown in Fig. 6-(b). Notably, Most of the negative samples are concentrated in the CoAP’s RFC 9177. This suggests that aiocoap has likely not yet implemented the specifications outlined in RFC 9177. This hypothesis aligns with the author’s introduction on GitHub (aiocoap, 2025), which specifies the standards supported by aiocoap, confirming that RFC 9177 has not yet been implemented in aiocoap. By generating as many positive samples as possible, iPanda can effectively reduce the scope of conformance testing. In addition, the test results of negative samples also guide further testing.
5. Related work
Conformance testing tools. Several tools have been developed for communication protocol conformance testing. Testing and Test Control Notation version 3 (TTCN-3) is a widely used language for rigorous conformance certification in mobile communications, and IoT protocols (Grabowski et al., 2003). Scapy enables flexible construction and transmission of custom packets via scripting, aiding tests of protocol edge cases and exception handling, particularly in security assessments (Rohith et al., 2018). Protocol fuzz testing tools, such as Fairfuzz and Boofuzz, have demonstrated strong capabilities in uncovering vulnerabilities and testing robustness (Pereyda, 2019; Lemieux and Sen, 2018; Peng et al., 2018). However, these tools are typically applied at specific stages within existing testing workflows and heuristic-based approaches (Makhmudov et al., 2025). A major limitation remains their dependence on the manual creation of test cases and scripts, requiring substantial developer effort. Currently, no tool seamlessly integrates protocol documentation, implementation, and tests into an end-to-end conformance testing process. Addressing this gap is precisely the focus of our work.
LLM-Based automated testing tools. Automated testing tools based on LLMs have gained increasing attention in recent years. Researchers have begun exploring LLM-powered agents for automating testing tasks. In code testing, some studies leveraged LLMs to generate high-coverage unit tests, such as those for the JUnit testing framework (Siddiq et al., 2024; Guilherme and Vincenzi, 2023). LLIFT (Li et al., 2023) employs LLMs for static code analysis to detect potential security vulnerabilities. PENTESTGPT, an LLM-based penetration testing tool, effectively identifies common vulnerabilities and analyzes source code for flaws (Deng et al., 2023b). DB-GPT integrates the Tree-of-Thought method into LLMs to systematically analyze database anomalies (Zhou et al., 2023). In fuzz testing, existing tools often struggle with message format obfuscation and dependencies between messages. To address this, LLMIF (Wang et al., 2024) integrates LLMs into IoT fuzz testing to automate the extraction of protocol formats and device response inference. However, current LLM-driven testing tools typically target isolated tasks and do not effectively support conformance testing. To address this research gap, we introduce iPanda, an LLM-based agent tailored specifically for protocol conformance testing.
Augmented LLMs. Although LLMs excel in tasks such as question answering and text generation, their performance remains constrained by the limitations of their pre-trained datasets and the contextual information provided during inference. As a result, they may perform poorly in certain specialized tasks. To address these limitations, researchers have explored various tools to enhance LLM capabilities, including Web browsers (Deng et al., 2023a; Nakano et al., 2021; Chen et al., [n. d.]), RAG (Du et al., 2024; Jeong, 2023; Ng et al., 2025), programming tools (Alinezhad et al., 2019; Lu et al., 2023), other deep neural network models (Shen et al., 2023), and so on (Wang et al., 2022; Yao et al., 2023; Shinn et al., 2023). Similarly, in our work, iPanda incorporates RAG and programming tools to enhance the effectiveness of LLMs, improving their ability to analyze protocol specifications, generate test cases, and interact with protocol implementation libraries.
6. Conclusion
We present iPanda, an LLM-based intelligent agent for automated, end-to-end conformance testing of communication protocols. iPanda leverages the keyword-based TCG method to efficiently generate test cases from protocol documents, and employs code-oriented prompting methods (e.g., codeRAG and Augmented Role Prompting) to enhance its understanding of tested protocol implementation and generate test code. Furthermore, the augmented CRITIC is integrated for iterative code refinement. Through execution-based validation, iPanda assesses whether the protocol implementation adheres to the specified protocol requirements. Experimental results prove the effectiveness and high efficiency of iPanda in conformance testing.